Margin, Turnover, ROI

Required:

1. Consider the data for each of the following four independent companies. Calculate the missing values in the table below. For margin and ROI, enter your answers as percentages, rounded to two decimal places. For example, the decimal value .03827 would be entered as "3.83" percent. For turnover, enter your answer as a decimal value rounded to two decimal places.

   A    B    C    D
Revenue $10,000 $49,000 $96,000 $
Expenses $8,000 $ $90,240 $
Operating income $2,000 $12,250 $ $
Assets $40,000 $ $48,000 $9,500
Margin % 25 % % 6.00 %
Turnover 0.50 2.00
ROI % % % %

2. Assume that the cost of capital is 9 percent for each of the four firms. Compute the residual income for each of the four firms. If the residual income is negative, enter a negative amount.

A's residual income $
B's residual income $
C's residual income $
D's residual income $
Answer #1

1. Consider the data for each of the following four independent companies. Calculate the missing values in the table below. For margin and ROI, enter your answers as percentages, rounded to two decimal places. For example, the decimal value .03827 would be entered as "3.83" percent. For turnover, enter your answer as a decimal value rounded to two decimal places.

   A    B    C    D
Revenue $10,000 $49,000 $96,000 1140*100/6 = 19000
Expenses $8,000 49000-12250 = 36750 $90,240 $19000-1140 = 17860
Operating income $2,000 $12,250 $96000-90240 = 5760 9500*12% = 1140
Assets $40,000 49000*2 = 98000 $48,000 $9,500
Margin 2000/10000 = 20 % 25 % 5760/96000 = 6% % 6.00 %
Turnover 10000/40000 = 0.25 0.50 96000/48000 = 2 2.00
ROI 2000/40000 = 5 % 12250/98000 = 12.5 % 5760/48000 = 12% % 12 %

2. Assume that the cost of capital is 9 percent for each of the four firms. Compute the residual income for each of the four firms. If the residual income is negative, enter a negative amount.

A's residual income 2000-(40000*9%) = -1600
B's residual income 12250-(98000*9%) = 3430
C's residual income 5760-(48000*9%) = 1440
D's residual income 1140-(9500*9%) = 285